OpenStudy (anonymous):

Solve x^2 + 4x + 8 = 0

OpenStudy (anonymous):

do you know the quadratic formula ? Discriminant ?

OpenStudy (anonymous):

I got -4 plus or negative 8i but it's not in the choices so i did something wrong. If you can factor them are you supposed to?

OpenStudy (anonymous):

(x+2)^2=-4

OpenStudy (anonymous):

x=-2+-2i

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Discriminant=b^2-4ac=16-4*8=16-32=-16